OpenStudy (anonymous):

Solve the inequality -x^3-4x

OpenStudy (anonymous):

\[0\leq x^3+4x\] \[0\leq x(x^2+4)\]and now since \(x^2+4\geq 4\) only need \(0\leq x\)

OpenStudy (anonymous):

So the answer is 0<=x

OpenStudy (anonymous):

that is what i get, yes